# MigGPT: Harnessing Large Language Models for Automated Migration of Out-of-Tree Linux Kernel Patches Across Versions

### v3

@@ 1 Introduction
With the substantial progress made by Large Language Models (LLMs) in understanding rf-1 ; rf-2 ; rf-3 and generating code codebert ; rf-4 ; rf-5 ; intro-d-1 ; intro-d-2 , there is a promising opportunity to leverage LLMs for the automated migration and maintenance of out-of-tree kernel patches. However, due to the inherent lack of determinism in LLMs when generating content, several challenges arise when directly employing these models to handle the migration and maintenance of out-of-tree kernel patches. These challenges include 1) structural similarity-induced failure, 2) non-deterministic code snippet boundaries, 3) absence of associated code snippet information, and 4) inaccuracies in locating migration points, which reveals that LLMs struggle with incomplete code context understanding and inaccurate migration point identification.
@@ 5.4 Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3 encountered during the migration update of out-of-tree kernel patches by LLMs. In challenge 1, LLMs are prone to be misled by similar function structure when processing function definitions in code snippets, leading to incorrect retrieval of $s_{\text{new}}$ in $\text{file}_{\text{new}}$, which ultimately results in erroneous migrated $s^{\prime}_{\text{new}}$. To overcome this challenge, it is necessary to constrain the LLM’s attention to the target code snippet. As illustrated in Figure 2, the retrieval augmentation module achieves this by constructing a code fingerprint structure ($\text{CFP}_{\text{old}}$) for the old version of the Linux kernel code snippets $s_{\text{old}}$. By analyzing $\text{CFP}_{\text{old}}$, the module extracts the function signatures of the function definitions contained within $s_{\text{old}}$. These function signatures are then used to build a prompt to describe the structure information (“Structure Prompt"), which is incorporated into the input fed to the LLM. An example is provided in the App. D.2.
@@ 5.5 Retrieval Alignment Module
The retrieval alignment module is devised to tackle challenge 2, which was encountered during the migration update of out-of-tree kernel patches by LLMs. Challenge 2 indicates that when the LLM retrieves the target code snippet $s_{\text{new}}$ from the new version of the Linux kernel file $\text{file}_{\text{new}}$, there can be a mismatch between the boundary line of $s_{\text{new}}$. To address this issue, we need to leverage the information from the first and last lines of the old version code snippet $s_{\text{old}}$ to aid in the localization during the retrieval of $s_{\text{new}}$. As illustrated in Figure 2, we utilize the code fingerprint structure $\text{CFP}_{\text{old}}$ of $s_{\text{old}}$. By taking advantage of its linear structure, we obtain the CFP statements for the first and last lines. These statements are used to construct an “Alignment Prompt", which describes the information of the first and last lines and is included as part of the input to the LLM. This prompt guides the LLM in performing the retrieval task better by accurately identifying the boundaries of the code snippet.
@@ 6.1 Evaluation Settings
We assess MigGPT using two benchmarks: the out-of-tree kernel patch migration benchmark from Section 4 and FixMorph’s CVE patch backporting benchmark FixMorph , which includes 350 instances. For baselines, we use vanilla LLMs, including GPT-3.5 gpt-3.5 , GPT-4-turbo gpt4 , OpenAI-o1 openai-o1 , DeepSeek-V2.5 deepseek-2.5 , DeepSeek-V3 deepseek-3 , Deepseek-R1 deepseek-r1 , Llama-3.1-8B and Llama-3.1-70B-Instruct llama-3.1 , as they are widely recognized for their advanced performance, along with previous migration efforts like FixMorph FixMorph , TSBPORT TSBPORT and PPatHF PPatHF . Evaluation metrics include “best match" (exact code similarity after removing spaces, line breaks, and tab characters), “semantic match" (CodeBLEU with a 0.9 threshold for binary classification, detailed in App. F.6codebleu , and “human match" (developer-judged functional equivalence, detailed in App. F.3). We also considered the compilation success rate, as detailed in App. F.4. The hyperparameter $m$ is set to 3.
@@
Figure 5: The semantic match accuracy of target code snippets retrieval task and target code snippets migration task across various LLMs. “One-step" indicates the direct utilization of an LLM to complete the migration task in a single step.
@@ 6.5 Failure Analysis (RQ3)
In our in-depth analysis of "human match" errors in MigGPT-augmented GPT-4-turbo, we identified the following primary categories:
@@ 7 Conclusion
This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grants No.XDB0660300, XDB0660301, XDB0660302), Science and Technology Major Special Program of Jiangsu (Grant No. BG2024028), the NSF of China (Grants No. U22A2028, 62302483, 6240073476), CAS Project for Young Scientists in Basic Research (YSBR-029) and Youth Innovation Promotion Association CAS.
@@ A.1 Collection
The migration examples in our benchmark are derived from three open-source out-of-tree kernel patch projects: RT-PREEMPT [21], HAOC [16] and Raspberry Pi kernel [39]. RT-PREEMPT  [^3] https://mirrors.edge.kernel.org/pub/linux/kernel/projects/rt enhances the Linux kernel’s real-time performance for timing-sensitive applications like industrial control and robotics, while Raspberry Pi Linux  [^4] https://github.com/raspberrypi/linux offers a lightweight kernel optimized for embedded systems. HAOC  [^5] https://gitee.com/src-openeuler/kernel/blob/master/ improves kernel security through a "dual-kernel" design, enhancing code behavior, data access, and permission management. These projects are widely adopted in industry and open-source communities, ensuring coverage of critical out-of-tree patch scenarios. Notably, RT-PREEMPT’s latest version has been integrated into the mainline Linux kernel for maintenance and no longer exists as an out-of-tree kernel patch. However, this does not impede our utilization of it for research on automated migration and maintenance of out-of-tree kernel patches.
@@ C.4 Challenge 4
As shown in Figure 11, to accurately obtain the migrated out-of-tree kernel patch code $s^{\prime}_{\text{new}}$, it is essential to perform two modifications on the new Linux kernel code segment $s_{\text{new}}$ (specifically, adding the code snippet #ifdef CONFIG_HIVE at two locations). However, when LLMs undertake this task, they either misidentify the migration positions or only execute one of the required modifications. This results in the failure of the out-of-tree kernel patch code migration. This issue reveals the limitations of LLMs in interpreting the precise context required for accurate migration, suggesting a need for more refined techniques to enhance the models’ ability to infer migration points based on the given information correctly.
@@ D.2 Examples of Retrieval Augmentation Module
For challenge 1, we construct a “Structure Prompt" to specify the signatures of the code snippet $s_{\text{old}}$. By constructing the code fingerprint structure $\text{CFP}_{\text{old}}$ from $s_{\text{old}}$ as shown in Figure 8, we can extract FuncDef statements that contain the code signatures (Figure 13), thereby generating a “Structure Prompt" that describes these signatures. Consequently, the LLM will focus its attention on the function definition __pte_free_tlb rather than on the similar function definition __pmd_free_tlb. This Structure prompt enhances the LLM’s ability by providing a precise description of the target code, allowing the LLM to focus more accurately on the relevant code snippet and improving the precision of the retrieval.
@@ D.3 Examples of Migration Augmentation Module
The migration augmentation module is primarily designed to address challenge 4. Specifically, as shown in Figure 11, we conduct a comparative analysis between the code fingerprint structures $\text{CFP}_{\text{old}}$ and $\text{CFP}^{\prime}_{\text{old}}$ of the code snippets to ascertain that there are two primary migration points. The first point is located after the comment statement Tial call offset... and before the macro definition statement #define PROLOGUE_OFFSET.... The second point is situated after the statement const struct bpf_prog... and before the statement const int idx0=ctx->idx. By constructing the “Location Prompt", we enable the LLM to precisely locate the migration points, thereby successfully completing the task of migrating and maintaining the out-of-tree kernel patch.
@@
11:+RetrievalTaskPrompt+StructurePrompt+RetrievalTaskPrompt+StructurePrompt
@@ Appendix E More Results of MigGPT
We have also tested the performance of MigGPT on DeepSeek-V2.5 [7], as shown in Table 11 and Table 12, MigGPT outperforms the vanilla LLM.
@@ F.6 Threshold of CodeBLEU
CodeBLEU [40] is an automated metric designed to evaluate the quality of code generation, specifically tailored for tasks involving the generation of programming code. By integrating both syntactic and semantic features of code, CodeBLEU provides a similarity score ($[0,1]$) between two code snippets. We employ CodeBLEU as a measure of “semantic match” and investigate the alignment between CodeBLEU-based “semantic matches” and “human matches” across various thresholds. As illustrated in Figure 15 and Table 15, we identify a threshold of 0.9 as optimal for our proposed benchmark, ensuring a high degree of consistency between “semantic matches” derived from CodeBLEU and those determined by human evaluation.
@@ G.1 Type 1 Migration Sample
Traditional methods like FixMorph [43], SyDIT [31], TSBPORT [53], and PPatHF [37] target single-step migration and cannot solve the target code retrieval problem, while MigGPT tackles both problems, which is harder and more practical. Besides, traditional methods struggle with Type 1 cases due to their reliance on static alignment and predefined transformation rules. As an example shown in Figure 17, when backporting a patch that modifies compute_stats() in the old kernel, FixMorph relies on static alignment (e.g., matching function names like process_data). However, since compute_stats() is now a standalone function in the new kernel, traditional methods like FixMorph cannot generate transformation rules for the split code structure, as its AST differencing assumes code blocks stay within the same function. This illustrates how traditional methods struggle with Type 1’s non-conflicting but structurally divergent changes.
@@ G.2 CFP and Intermediate Representations
  • Non-Compilable Code Support: Experimental/unmerged patches (e.g., ARM64-specific optimizations) may fail compilation, making IR extraction impossible. CFP operates directly on source fragments, even for "broken" code in development.
@@ G.3 Generalization of MigGPT
MigGPT demonstrates innovative advancements through its dual strengths of generalizability and domain-specific optimization. While initially designed for out-of-tree kernel patches, its core methodology addresses universal challenges in LLM-based code migration, such as resolving structure conflicts, precisely aligning code boundaries, reconstructing missing context, and establishing migration localization mechanisms, forming a framework extensible to multiple domains (CVE patch backporting [43, 53], forked code porting [37, 55]). The architecture incorporates its Code Fingerprint structure, which embeds Linux kernel-specific optimizations through preserving inline assembly instructions and kernel macro patterns, interpreting kernel-specific comment conventions, and adapting to coding norms like function chaining. This technical design achieves a balance between cross-domain adaptability and deep specialization, with the CFP serving as a modular component that enhances Linux kernel patch migration while maintaining the system’s capacity for expansion into other technical ecosystems. The solution thus enables bidirectional scalability, supporting both horizontal domain transfer and vertical technical refinement.
@@ G.4 Structured Analysis of CFP
The high-level idea of using program analysis techniques to create structured representations of code to guide LLMs is an emerging area of research. We have compared MigGPT with existing research in this emerging area. For instance, works such as AutoOS [3] and BYOS [29] utilize heuristic tree structures to assist LLMs in optimizing kernel configurations during the deployment of operating systems. Similarly, LLM-based fuzzing tools [49, 18] employ formal templates to guide LLMs in generating more effective test cases. These approaches, much like MigGPT, leverage deterministic information such as trees, graphs, and lists to mitigate the inherent uncertainty and randomness of LLMs, thereby improving performance across various application tasks. However, it’s worth noting that the specific deterministic information used and its data structure are often tailored to the particular task at hand.

### v4

@@ 1 Introduction
With the substantial progress made by Large Language Models (LLMs) in understanding rf-1; rf-2; rf-3 and generating code codebert; rf-4; rf-5; intro-d-1; intro-d-2, there is a promising opportunity to leverage LLMs for the automated migration and maintenance of out-of-tree kernel patches. However, due to the inherent lack of determinism in LLMs when generating content, several challenges arise when directly employing these models to handle the migration and maintenance of out-of-tree kernel patches. These challenges include 1) structural similarity-induced failure, 2) non-deterministic code snippet boundaries, 3) absence of associated code snippet information, and 4) inaccuracies in locating migration points, which reveal that LLMs struggle with incomplete code context understanding and inaccurate migration point identification.
@@ 5.4 Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3 encountered during the migration update of out-of-tree kernel patches by LLMs. In challenge 1, LLMs are prone to be misled by similar function structure when processing function definitions in code snippets, leading to incorrect retrieval of $s_{\text{new}}$ in $\text{file}_{\text{new}}$, which ultimately results in erroneous migrated $s^{\prime}_{\text{new}}$. To overcome this challenge, it is necessary to constrain the LLM’s attention to the target code snippet. As illustrated in Figure 2, the retrieval augmentation module achieves this by constructing a code fingerprint structure ($\text{CFP}_{\text{old}}$) for the old version of the Linux kernel code snippets $s_{\text{old}}$. By analyzing $\text{CFP}_{\text{old}}$, the module extracts the function signatures of the function definitions contained within $s_{\text{old}}$. These function signatures are then used to build a prompt to describe the structure information (“Structure Prompt”), which is incorporated into the input fed to the LLM. An example is provided in the App. D.2.
@@ 5.5 Retrieval Alignment Module
The retrieval alignment module is devised to tackle challenge 2, which was encountered during the migration update of out-of-tree kernel patches by LLMs. Challenge 2 indicates that when the LLM retrieves the target code snippet $s_{\text{new}}$ from the new version of the Linux kernel file $\text{file}_{\text{new}}$, there can be a mismatch between the boundary line of $s_{\text{new}}$. To address this issue, we need to leverage the information from the first and last lines of the old version code snippet $s_{\text{old}}$ to aid in the localization during the retrieval of $s_{\text{new}}$. As illustrated in Figure 2, we utilize the code fingerprint structure $\text{CFP}_{\text{old}}$ of $s_{\text{old}}$. By taking advantage of its linear structure, we obtain the CFP statements for the first and last lines. These statements are used to construct an “Alignment Prompt”, which describes the information of the first and last lines and is included as part of the input to the LLM. This prompt guides the LLM in performing the retrieval task better by accurately identifying the boundaries of the code snippet.
@@ 6.1 Evaluation Settings
We assess MigGPT using two benchmarks: the out-of-tree kernel patch migration benchmark from Section 4 and FixMorph’s CVE patch backporting benchmark FixMorph, which includes 350 instances. For baselines, we use vanilla LLMs, including GPT-3.5 gpt-3.5, GPT-4-turbo gpt4, OpenAI-o1 openai-o1, DeepSeek-V2.5 deepseek-2.5, DeepSeek-V3 deepseek-3, Deepseek-R1 deepseek-r1, Llama-3.1-8B and Llama-3.1-70B-Instruct llama-3.1, as they are widely recognized for their advanced performance, along with previous migration efforts like FixMorph FixMorph, TSBPORT TSBPORT and PPatHF PPatHF. Evaluation metrics include “best match” (exact code similarity after removing spaces, line breaks, and tab characters), “semantic match” (CodeBLEU with a 0.9 threshold for binary classification, detailed in App. F.6codebleu, and “human match” (developer-judged functional equivalence, detailed in App. F.3). We also considered the compilation success rate, as detailed in App. F.4. The hyperparameter $m$ is set to 3.
@@
Figure 5: The semantic match accuracy of the target code snippets retrieval task and the target code snippets migration task across various LLMs. “One-step” indicates the direct utilization of an LLM to complete the migration task in a single step.
@@ 6.5 Failure Analysis (RQ3)
In our in-depth analysis of “human match” errors in MigGPT-augmented GPT-4-turbo, we identified the following primary categories:
@@ 7 Conclusion
This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grants No.XDB0660300, XDB0660301, XDB0660302), Science and Technology Major Special Program of Jiangsu (Grant No. BG2024028), the NSF of China (Grants No. U22A2028, 62302483, 6240073476), CAS Project for Young Scientists in Basic Research (YSBR-029) and Youth Innovation Promotion Association CAS. This work is also supported by NSFC-92467102.
@@ A.1 Collection
The migration examples in our benchmark are derived from three open-source out-of-tree kernel patch projects: RT-PREEMPT [rt-preempt], HAOC [haoc-patch], and Raspberry Pi kernel [raspberrypi-linux]. RT-PREEMPT  [^3] https://mirrors.edge.kernel.org/pub/linux/kernel/projects/rt enhances the Linux kernel’s real-time performance for timing-sensitive applications like industrial control and robotics, while Raspberry Pi Linux  [^4] https://github.com/raspberrypi/linux offers a lightweight kernel optimized for embedded systems. HAOC  [^5] https://gitee.com/src-openeuler/kernel/blob/master/ improves kernel security through a “dual-kernel” design, enhancing code behavior, data access, and permission management. These projects are widely adopted in industry and open-source communities, ensuring coverage of critical out-of-tree patch scenarios. Notably, RT-PREEMPT’s latest version has been integrated into the mainline Linux kernel for maintenance and no longer exists as an out-of-tree kernel patch. However, this does not impede our utilization of it for research on automated migration and maintenance of out-of-tree kernel patches.
@@ C.4 Challenge 4
As shown in Figure 11, to accurately obtain the migrated out-of-tree kernel patch code $s^{\prime}_{\text{new}}$, it is essential to perform two modifications on the new Linux kernel code segment $s_{\text{new}}$ (specifically, adding the code snippet #ifdef CONFIG_HIVE at two locations). However, when LLMs undertake this task, they either misidentify the migration positions or only execute one of the required modifications. This results in the failure of the out-of-tree kernel patch code migration. This issue reveals the limitations of LLMs in interpreting the precise context required for accurate migration, suggesting a need for more refined techniques to enhance the models’ ability to correctly infer migration points based on the given information.
@@ D.2 Examples of Retrieval Augmentation Module
For challenge 1, we construct a “Structure Prompt” to specify the signatures of the code snippet $s_{\text{old}}$. By constructing the code fingerprint structure $\text{CFP}_{\text{old}}$ from $s_{\text{old}}$ as shown in Figure 8, we can extract FuncDef statements that contain the code signatures (Figure 13), thereby generating a “Structure Prompt” that describes these signatures. Consequently, the LLM will focus its attention on the function definition __pte_free_tlb rather than on the similar function definition __pmd_free_tlb. This Structure prompt enhances the LLM’s ability by providing a precise description of the target code, allowing the LLM to focus more accurately on the relevant code snippet and improving the precision of the retrieval.
@@ D.3 Examples of Migration Augmentation Module
The migration augmentation module is primarily designed to address challenge 4. Specifically, as shown in Figure 11, we conduct a comparative analysis between the code fingerprint structures $\text{CFP}_{\text{old}}$ and $\text{CFP}^{\prime}_{\text{old}}$ of the code snippets to ascertain that there are two primary migration points. The first point is located after the comment statement Tial call offset... and before the macro definition statement #define PROLOGUE_OFFSET.... The second point is situated after the statement const struct bpf_prog... and before the statement const int idx0=ctx->idx. By constructing the “Location Prompt”, we enable the LLM to precisely locate the migration points, thereby successfully completing the task of migrating and maintaining the out-of-tree kernel patch.
@@
11:$RetrievalTaskPrompt+StructurePrompt$
@@ Appendix E More Results of MigGPT
We have also tested the performance of MigGPT on DeepSeek-V2.5 [deepseek-2.5], as shown in Table 11 and Table 12, MigGPT outperforms the vanilla LLM.
@@ F.6 Threshold of CodeBLEU
CodeBLEU [codebleu] is an automated metric designed to evaluate the quality of code generation, specifically tailored for tasks involving the generation of programming code. By integrating both syntactic and semantic features of code, CodeBLEU provides a similarity score ($[0,1]$) between two code snippets. We employ CodeBLEU as a measure of “semantic match” and investigate the alignment between CodeBLEU-based “semantic matches” and “human matches” across various thresholds. As illustrated in Figure 15 and Table 15, we identify a threshold of 0.9 as optimal for our proposed benchmark, ensuring a high degree of consistency between “semantic matches” derived from CodeBLEU and those determined by human evaluation.
@@ G.1 Type 1 Migration Sample
Traditional methods like FixMorph [FixMorph], SyDIT [sydit], TSBPORT [TSBPORT], and PPatHF [PPatHF] target single-step migration and cannot solve the target code retrieval problem, while MigGPT tackles both problems, which is harder and more practical. Besides, traditional methods struggle with Type 1 cases due to their reliance on static alignment and predefined transformation rules. As an example shown in Figure 17, when backporting a patch that modifies compute_stats() in the old kernel, FixMorph relies on static alignment (e.g., matching function names like process_data). However, since compute_stats() is now a standalone function in the new kernel, traditional methods like FixMorph cannot generate transformation rules for the split code structure, as its AST differencing assumes code blocks stay within the same function. This illustrates how traditional methods struggle with Type 1’s non-conflicting but structurally divergent changes.
@@ G.2 CFP and Intermediate Representations
  • Non-Compilable Code Support: Experimental/unmerged patches (e.g., ARM64-specific optimizations) may fail compilation, making IR extraction impossible. CFP operates directly on source fragments, even for“broken” code in development.
@@ G.3 Generalization of MigGPT
MigGPT demonstrates innovative advancements through its dual strengths of generalizability and domain-specific optimization. While initially designed for out-of-tree kernel patches, its core methodology addresses universal challenges in LLM-based code migration, such as resolving structure conflicts, precisely aligning code boundaries, reconstructing missing context, and establishing migration localization mechanisms, forming a framework extensible to multiple domains (CVE patch backporting [FixMorph, TSBPORT], forked code porting [PPatHF, add1]). The architecture incorporates its Code Fingerprint structure, which embeds Linux kernel-specific optimizations through preserving inline assembly instructions and kernel macro patterns, interpreting kernel-specific comment conventions, and adapting to coding norms like function chaining. This technical design achieves a balance between cross-domain adaptability and deep specialization, with the CFP serving as a modular component that enhances Linux kernel patch migration while maintaining the system’s capacity for expansion into other technical ecosystems. The solution thus enables bidirectional scalability, supporting both horizontal domain transfer and vertical technical refinement.
@@ G.4 Structured Analysis of CFP
The high-level idea of using program analysis techniques to create structured representations of code to guide LLMs is an emerging area of research. We have compared MigGPT with existing research in this emerging area. For instance, works such as AutoOS [r-1] and BYOS [r-2] utilize heuristic tree structures to assist LLMs in optimizing kernel configurations during the deployment of operating systems. Similarly, LLM-based fuzzing tools [r-3, r-4] employ formal templates to guide LLMs in generating more effective test cases. These approaches, much like MigGPT, leverage deterministic information such as trees, graphs, and lists to mitigate the inherent uncertainty and randomness of LLMs, thereby improving performance across various application tasks. However, it’s worth noting that the specific deterministic information used and its data structure are often tailored to the particular task at hand.